# Range Image-Based Implicit Neural Compression for LiDAR Point Clouds

### v1

@@ I Introduction
Light Detection and Ranging (LiDAR) sensors have gained significant attention not only in online applications but also in offline applications. In such offline applications, memory-efficient and precise three-dimensional (3D) scenes should be stored in advance and the 3D scenes should be smoothly retrieved from the storage based on the user demand for applications of 3D scene understanding such as digital archiving, environmental monitoring, navigation, and geological surveying  [1, 2, 3]. LiDAR sensors scan the physical space with the ego-centric coordinate and measure the distance to the closest point on surrounding objects for each angle, allowing the creation of a point cloud with 3D points corresponding to the intersection of laser beams with objects ahead. As the resolution of LiDAR sensors increases, effectively storing and transmitting LiDAR scans becomes a significant challenge, primarily due to the substantial volume of each LiDAR sequence.
@@
Although LiDAR scans are typically represented as 3D point clouds, they can also be expressed as a single-channel image, referred to as a two-dimensional (2D) range image (RI) [4], where point clouds captured in an ego-centric spherical coordinate system are projected onto a panoramic image. The x-y axes of the RI image correspond to the azimuth and elevation angles in the 3D spherical coordinate system, while each pixel value represents the distance to the corresponding point in that direction. Whereas 3D point clouds require $3N$ values to represent the locations of $N$ point measurements, RI s require only $N$ pixels at most, thus demonstrating their compactness.
@@
We can pursue methods to further compress RI s. One potential solution is to adapt conventional lossy image compression techniques, such as Joint Photographic Experts Group (JPEG) [5] and Joint Photographic Experts Group 2000 (JPEG2000) [6]. However, these methods are based on integer precision for pixel values, which is incompatible for RI s whose pixels are represented by single or double precision floating-point numbers to precisely express the distance to the corresponding point measurements. We can still utilize these compression methods by adjusting the bit precision of RI s to align with them, but this naturally leads to degradation in 3D point cloud reconstruction performance due to inadequate distance resolution. Another drawback of conventional compression methods lies in their strategy: they use block-based discrete cosine transform (DCT) and apply coarse quantization to high-frequency components based on human visual perception. This approach leads to significant degradation in the decoding performance of RI s, as RI s are characterized by large and sudden changes in the pixel value between foreground objects and distant backgrounds or pixels without any assigned point measurement.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
To effectively compress RI s while preserving their fine precision and high-frequency changes in pixel value, this paper presents a novel RI compression method inspired by implicit neural representation (INR)-based image compression technique [7]. INR [8, 9] is a lightweight representation of multidimensional signals by compressing them into shallow neural networks (NNs). Specifically, INR overfits NNs with a limited number of parameters to the signals of interest through supervised learning, and the trained parameters become the compressed signal representation by providing the mapping function from signal indices, for example, coordinates on the image plane, to the corresponding signals. A primary challenge for INR-based signal compression is to ensure the precision of high-frequency details while simultaneously managing the constraints imposed by the limited model size. To address this challenge, we propose an extended INR training approach that incorporates both a mask image and an RI for point depth information. The mask image is a binary map that indicates whether each pixel corresponds to a projected 3D point (zero) or not (one). We begin by generating a mask image from the RI, followed by learning two separate coordinate-to-value mappings using distinct INR architectures: one for depth INR and one for the mask INR. During decoding, these trained INR s reconstruct both the depth and mask images, and the final reconstructed RI is obtained by applying the mask to the depth image. Although the reconstructed depth image may contain values for pixels that do not correspond to any 3D points, applying the mask enforces hard thresholding, effectively removing these artifacts. This process ensures a high-quality reconstructed RI, with sharp edges accurately preserved.
@@
  • We evaluate our proposed method using KITTI dataset [10] and compare its performance with existing baselines including conventional image compression, point cloud compression (PCC) [11], and RI-based and INR-based image compression, and show the better rate-distortion (R-D) performance and downstream task quality of our method than the baselines.
@@ II-A Point Cloud Compression
The measured distance from LiDAR sensors is usually represented as a 3D point cloud. Each point cloud consists of a set of 3D points, and each point is defined by 3D coordinates, i.e., (X, Y, Z). The graph-based and tree-based compression methods have been proposed to compress coordinate information, that is, geometry information. The graph-based methods regard the 3D points as graph signals and define graph Fourier transform (GFT) for frequency conversion in the graph domain. [12, 13, 14] utilized GFT for the geometry compression. Other studies [15, 16] reduce the storage and transmission costs for graph signal reconstruction. The tree-based compression method is another popular strategy for compressing geometry information. The typical way is octree-based representation, such as point cloud library (PCL) and geometry-based point cloud compression (G-PCC) [17, 11]. Some recent studies have been proposed to improve the efficiency of geometry compression using traditional signal processing [18] and deep neural network (DNN) [19, 20] solutions, respectively. For example, the study in [18] adaptively adopts the quad-tree (QT) and binary-tree (BT) block partitions in addition to those of octrees to improve the efficiency of the coding.
@@ II-B LiDAR Range Image Compression
Many recent works consider projecting the measured LiDAR information onto 2D RI to represent the measured distance information in a compact format. There are two types of input LiDAR information to obtain the corresponding RI s: 1) raw packet containing the LiDAR laser IDs [21], the rotation angle of the LiDAR sensors and the distance values, and 2) 3D point clouds [22, 23, 24]. Our study utilizes 3D point clouds. The obtained RI s are then intra-coded [22] or inter-coded [23, 24] in lossless and lossy manners. Here, intra-coding reduces the spatial redundancy in each RI, whereas inter-coding reduces the temporal redundancy across RI s. In R-PCC [22], which is an intra-coding method, each RI can be coded by using a lossless coding method, such as LZ4 and Deflate, to compress the floating-point format. Our study is designed for RI intra-coding and exploits the INR-based compression to represent LiDAR measurements in small storage and transmission costs.
(a) Patched depth image and mask image from LiDAR point cloud
(b) Proposed encoder and decoder
@@ II-C Implicit Neural Compression
The existing INR-based compression can be classified into pixel-wise, patch-wise, and frame-wise architectures. The frame-wise architectures realized inter-coding between multiple video frames to remove temporal redundancy. They feed the frame index and/or the corresponding embeddings to the NNs to generate each frame. Neural Representations for Videos (NeRV) [25] is the first work on frame-wise video compression, and various extensions [26, 27, 28, 29, 30, 31, 32, 33, 34, 35] are proposed to improve the quality of reconstruction. However, NeRV architectures are large models when used for intra-coding each image.
The pixel-wise INR architecture [7, 36] takes the pixel index as input and reconstructs the corresponding pixel value. The patch-wise INR architecture was first proposed in [37]. Specifically, each image is divided into multiple patches, and the INR architecture takes the patch index as input to exploit the similarity of local adjacent pixels for high-quality reconstruction under the same model size. A key issue in such INR architectures is the lack of precision in high-frequency details with a small NN architecture. To represent high-frequency details under a small NN architecture, SIREN in [8] argues that sinusoidal activations work better than Rectified Linear Unit (ReLU) networks because sinusoidal activations can fit signals contained in higher-order derivatives. To address the same problem, our paper extends the training process of INRs by separating RI s into depth and mask images.
@@ III-A Overview
Fig. 1 shows an end-to-end architecture of the proposed scheme. Fig. 1 (a) specifically shows the procedure to obtain RI and corresponding depth and mask images from the LiDAR 3D point cloud. We consider that the LiDAR measurement to be compressed is a 3D point cloud consisting of $N$ points, denoted as $\mathbf{P}=\{\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]\ |\ i=1,\cdots,N\}$, where $x_{i},y_{i},z_{i}\in\mathbb{R}$ represent the Cartesian coordinates of the $i$-th point. The point cloud is first transformed into the spherical coordinate system. Subsequently, each point is projected onto a 2D image plane by mapping it to a pixel in a single-channel image $I\in\mathbb{R}^{W\times H}$, producing an RI. The RI is then divided into a depth image $I_{D}\in\mathbb{R}^{W\times H}$ and a mask image $I_{M}\in\{0,1\}^{W\times H}$. The depth image $I_{D}$ is further segmented into small rectangular regions, or patches, with a resolution of $\frac{W}{N_{p}}\times\frac{H}{N_{p}}$, where $N_{p}$ is the scaling factor.
Fig. 1 (b) shows the sequential operations of the encoder and decoder. In the encoder, two distinct INR s, namely the mask INR $\Phi(\cdot;\bm{\psi})$ and the depth INR $\Phi(\cdot;\bm{\omega})$ with learnable parameters $\bm{\psi}$ and $\bm{\omega}$, are trained to be overfitted to the depth image and the mask image, respectively. This training process is a pixel-wise process, which means that the parameters are trained to obtain a mapping from the pixel coordinates or patch indices on the images to their corresponding pixel values. This is achieved by sequentially providing a pair of indices to the networks. The well-trained parameters $\bm{\psi}$ and $\bm{\omega}$ are subsequently pruned and quantized as $\hat{\bm{\psi}}$ and $\hat{\bm{\omega}}$ to enhance their compactness, and we assume that these parameters are stored in storage or transmitted to content receivers as the lightweight format of the LiDAR measurements. In the decoding process, the decoder uses compressed parameters to reconstruct a mask image $\hat{I}_{M}$ and a depth image $\hat{I}_{D}$ from individual INR architectures $\Phi(\cdot;\hat{\bm{\psi}})$ and $\Phi(\cdot;\hat{\bm{\omega}})$, respectively. Similarly to the encoding process, the images are reconstructed by sequentially feeding a pair of coordinates and indices to the INR s and collecting all estimated values to form the shape of the image. We obtain the final result of RI, $\hat{I}$, by applying the mask image to the depth image to mask out any values of the pixels in the depth image corresponding to the pixels with a mask value of 1, indicating “no point”. Finally, the LiDAR point cloud is reconstructed as $\hat{\mathbf{P}}$ from $\hat{I}$ via a reverse coordinate projection process from the 2D image plane, through the spherical coordinate system, to the Cartesian coordinate system.
@@
[…]
@@ III-C Depth/Mask Image Construction
$I_{M}(u,v)=\begin{cases}1&\text{if }I(u,v)=\rho_{\rm null},\\ 0&\text{otherwise}.\end{cases}$ (3)
@@
$\mathcal{D}_{M}=\{((u,v),I_{M}(u,v))\mid$ $u\in\{1,\ldots,W\},v\in\{1,\ldots,H\}\}.$ (4)
@@
In addition, we divide the depth image into small rectangular areas, or patches, inspired by recent works [37] to improve the decoding performance and the quality of the reconstructed depth image. Specifically, the RI is evenly segmented into patches $I_{D}^{\prime}(i)\in\mathbb{R}^{\frac{W}{N_{p}}\times\frac{H}{N_{p}}}$, where $N_{p}$ is a scaling factor and $i=1,\cdots,N_{p}^{2}$. Each patch is assigned a patch index, allowing us to specify a pixel in the patched RI as $I_{D}^{\prime}(i,i_{u},i_{v})$, where $i$ represents the patch index and $i_{u},i_{v}$ denotes the in-patch pixel coordinates whose origin is the top left pixel in the $i$-th patch. Similarly to the mask image, we also construct a dataset $\mathcal{D}_{D}$ for training depth INR $\Psi(\cdot;\bm{\omega})$ which consists of pairs of a patch index, in-patch coordinates, and the corresponding depth values, excluding unassigned pixels, as follows:
$\mathcal{D}_{D}=\{((i,i_{u},i_{v}),I_{D}(i,i_{u},i_{v}))\mid$ $i\in\{1,\ldots,N_{p}^{2}\},$
@@ III-D INR-based RI Encoder
In the encoding process, the mask INR $\Phi(\cdot;\psi)$ and the depth INR $\Phi(\cdot;\omega)$ are trained to obtain good parameters to express the coordinate-to-value relationships contained in the mask dataset $\mathcal{D}_{M}$ and $\mathcal{D}_{D}$. Figs. 2 (a) and (b) show the detailed architecture of the proposed depth INR and mask INR, respectively.
@@ III-D2 Depth INR
Similar to the mask INR, we also assume the existence of another function $\Phi_{D}$ regarding the depth image, which maps the pair of a patch index and an in-patch pixel coordinate to a depth value as
$\Phi_{D}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{1}.$ (9)
and the objective of the depth INR is to obtain parameters $\bm{\omega}$ for good approximation of $\Phi_{D}$ as $\Phi((i,i_{u},i_{v});\bm{\omega})\approx\Phi_{D}$. Figs. 2 (b) shows the detailed architecture of the proposed depth INR. The depth INR shares its structure with the mask INR, with the exception of the input layer, which accepts 3 values. In the training process, pairs of patch index $i$ and an in-patch coordinate $(i_{u},i_{v})$ are sequentially passed to the depth INR network from the depth dataset $\mathcal{D}_{D}$, and the corresponding depth values are regressed. We employ mean squared error (MSE) loss as a regression loss function for the depth INR as
$\mathcal{L}_{\mathrm{MSE}}(\bm{\omega})=\frac{1}{HW}\sum_{i}^{N_{p}^{2}}\sum_{i_{u}}^{W/N_{p}}\sum_{i_{v}}^{H/N_{p}}\|\Phi((i,i_{u},i_{v});\bm{\omega})-I_{D}(i,i_{u},i_{v})\|^{2}.$ (10)

### v2

@@ I Introduction
LiDAR sensors have gained significant attention not only in online applications but also in offline applications. In such offline applications, memory-efficient and precise three-dimensional (3D) scenes should be stored in advance and the 3D scenes should be smoothly retrieved from the storage based on the user demand for applications of 3D scene understanding such as digital archiving, environmental monitoring, navigation, and geological surveying  [env_moniter, modeling, urban]. LiDAR sensors scan the physical space with the ego-centric coordinate and measure the distance to the closest point on surrounding objects for each angle, allowing the creation of a point cloud with 3D points corresponding to the intersection of laser beams with objects ahead. As the resolution of LiDAR sensors increases, effectively storing and transmitting LiDAR scans becomes a significant challenge, primarily due to the substantial volume of each LiDAR sequence.
@@
Although LiDAR scans are typically represented as 3D point clouds, they can also be expressed as a single-channel image, referred to as a two-dimensional (2D) range image (RI) [RI], where point clouds captured in an ego-centric spherical coordinate system are projected onto a panoramic image. The x-y axes of the RI image correspond to the azimuth and elevation angles in the 3D spherical coordinate system, while each pixel value represents the distance to the corresponding point in that direction. Whereas 3D point clouds require $3N$ values to represent the locations of $N$ point measurements, RI s require only $N$ pixels at most, thus demonstrating their compactness.
@@
We can pursue methods to further compress RI s. One potential solution is to adapt conventional lossy image compression techniques, such as Joint Photographic Experts Group (JPEG) [jpeg] and Joint Photographic Experts Group 2000 (JPEG2000) [jpeg2000]. However, these methods are based on integer precision for pixel values, which is incompatible for RI s whose pixels are represented by single or double precision floating-point numbers to precisely express the distance to the corresponding point measurements. We can still utilize these compression methods by adjusting the bit precision of RI s to align with them, but this naturally leads to degradation in 3D point cloud reconstruction performance due to inadequate distance resolution. Another drawback of conventional compression methods lies in their strategy: they use block-based discrete cosine transform (DCT) and apply coarse quantization to high-frequency components based on human visual perception. This approach leads to significant degradation in the decoding performance of RI s, as RI s are characterized by large and sudden changes in the pixel value between foreground objects and distant backgrounds or pixels without any assigned point measurement.
@@
To effectively compress RI s while preserving their fine precision and high-frequency changes in pixel value, this paper presents a novel RI compression method inspired by implicit neural representation (INR)-based image compression technique [dupont2021coin]. INR [sitzmann2020implicit, tancik2020fourier] is a lightweight representation of multidimensional signals by compressing them into shallow neural networks (NNs). Specifically, INR overfits NNs with a limited number of parameters to the signals of interest through supervised learning, and the trained parameters become the compressed signal representation by providing the mapping function from signal indices, for example, coordinates on the image plane, to the corresponding signals. A primary challenge for INR-based signal compression is to ensure the precision of high-frequency details while simultaneously managing the constraints imposed by the limited model size. To address this challenge, we propose an extended INR training approach that incorporates both a mask image and an RI for point depth information. The mask image is a binary map that indicates whether each pixel corresponds to a projected 3D point (one) or not (zero). We begin by generating a mask image from the RI, followed by learning two separate coordinate-to-value mappings using distinct INR architectures: one for depth INR and one for the mask INR. During decoding, these trained INR s reconstruct both the depth and mask images, and the final reconstructed RI is obtained by applying the mask to the depth image. Although the reconstructed depth image may contain values for pixels that do not correspond to any 3D points, applying the mask enforces hard thresholding, effectively removing these artifacts. This process ensures a high-quality reconstructed RI, with sharp edges accurately preserved.
@@
  • We evaluate our proposed method using KITTI dataset [kitti] and compare its performance with existing baselines including conventional image compression, point cloud compression (PCC) [bib:vpcc_gpcc], and RI-based and INR-based image compression, and show the better rate-distortion (R-D) performance and downstream task quality of our method than the baselines, especially at low bitrates.
@@ II-A Point Cloud Compression
The measured distance from LiDAR sensors is usually represented as a 3D point cloud. Each point cloud consists of a set of 3D points, and each point is defined by 3D coordinates, i.e., (X, Y, Z). The graph-based and tree-based compression methods have been proposed to compress coordinate information, that is, geometry information. The graph-based methods regard the 3D points as graph signals and define graph Fourier transform (GFT) for frequency conversion in the graph domain. [bib:graph_PCC2, bib:holoplus, bib:kirihara] utilized GFT for the geometry compression. Other studies [bib:ueno, bib:gsp_inr] reduce the storage and transmission costs for graph signal reconstruction. The tree-based compression method is another popular strategy for compressing geometry information. The typical way is octree-based representation, such as point cloud library (PCL) and geometry-based point cloud compression (G-PCC) [bib:PCL, bib:vpcc_gpcc]. Some recent studies have been proposed to improve the efficiency of geometry compression using traditional signal processing [bib:quad] and deep neural network (DNN) [OctAttention, huang2020octsqueeze] solutions, respectively. For example, the study in [bib:quad] adaptively adopts the quad-tree (QT) and binary-tree (BT) block partitions in addition to those of octrees to improve the efficiency of the coding.
@@ II-B LiDAR Range Image Compression
Many recent works consider projecting the measured LiDAR information onto 2D RI to represent the measured distance information in a compact format. There are two types of input LiDAR information to obtain the corresponding RI s: 1) raw packet containing the LiDAR laser IDs [riddle], the rotation angle of the LiDAR sensors and the distance values, and 2) 3D point clouds [wang2022r, bib:inter_rpcc, bib:segmentation]. Our study utilizes 3D point clouds. The obtained RI s are then intra-coded [wang2022r] or inter-coded [bib:inter_rpcc, bib:segmentation] in lossless and lossy manners. Here, intra-coding reduces the spatial redundancy in each RI, whereas inter-coding reduces the temporal redundancy across RI s. In R-PCC [wang2022r], which is an intra-coding method, each RI can be coded by using a lossless coding method, such as LZ4 and Deflate, to compress the floating-point format. Our study is designed for RI intra-coding and exploits the INR-based compression to represent LiDAR measurements in small storage and transmission costs.
((a))
((b))
@@ II-C Implicit Neural Compression
The existing INR-based compression can be classified into pixel-wise, patch-wise, and frame-wise architectures. The frame-wise architectures realized inter-coding between multiple video frames to remove temporal redundancy. They feed the frame index and/or the corresponding embeddings to the NNs to generate each frame. Neural Representations for Videos (NeRV) [chen2021nerv] is the first work on frame-wise video compression, and various extensions [hinerv, lee2023ffnerv, he2023dnerv, 2024VQNeRV, nirvana, NeRI, yan2024ds, Gomes_2023_CVPR, chen2023hnerv, 2024boosting] are proposed to improve the quality of reconstruction. However, NeRV architectures are large models when used for intra-coding each image.
The pixel-wise INR architecture [dupont2021coin, dupont2022coin++] takes the pixel index as input and reconstructs the corresponding pixel value. The patch-wise INR architecture was first proposed in [yunpengps-nerv]. Specifically, each image is divided into multiple patches, and the INR architecture takes the patch index as input to exploit the similarity of local adjacent pixels for high-quality reconstruction under the same model size. A key issue in such INR architectures is the lack of precision in high-frequency details with a small NN architecture. To represent high-frequency details under a small NN architecture, SIREN in [sitzmann2020implicit] argues that sinusoidal activations work better than Rectified Linear Unit (ReLU) networks because sinusoidal activations can fit signals contained in higher-order derivatives. To address the same problem, our paper extends the training process of INRs by separating RI s into depth and mask images.
@@ III-A Overview
Fig. 1 shows an end-to-end architecture of the proposed scheme. Fig. 1 (a) specifically shows the procedure to obtain RI and corresponding depth and mask images from the LiDAR 3D point cloud. We consider that the LiDAR measurement to be compressed is a 3D point cloud consisting of $N$ points, denoted as $\mathbf{P}=\{\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]\ |\ i=1,\cdots,N\}$, where $x_{i},y_{i},z_{i}\in\mathbb{R}$ represent the Cartesian coordinates of the $i$-th point. The point cloud is first transformed into the spherical coordinate system. Subsequently, each point is projected onto a 2D image plane by mapping it to a pixel in a single-channel image $I\in\mathbb{R}^{W\times H}$, producing an RI. Here, $W$ and $H$ represent the width and height of the range image, respectively. The RI is then divided into a depth image $I_{D}\in\mathbb{R}^{W\times H}$ and a mask image $I_{M}\in\{0,1\}^{W\times H}$. The depth image $I_{D}$ is further segmented into small rectangular regions, or patches, with a resolution of $\frac{W}{N_{p}}\times\frac{H}{N_{p}}$, where $N_{p}$ is the scaling factor.
[…]
@@
where $\phi_{\mathrm{up}}$ and $\phi_{\mathrm{down}}$ are the maximum and minimum value of $\phi$ in the dataset, $|\cdot|$ is the absolute value, and $\lfloor\cdot\rfloor$ is a floor function. $H$ and $W$ in Eq. (2) are the height and width of RI, and they are determined by the angular resolution of the LiDAR sensor for the elevation and azimuth axes. In this study, we set $H=64$ and $W=1024$. The value of each pixel $I(u,v)$ on the RI is the measured distance $\rho$, derived in Eq. (1), with arbitrary unit for physical length.
@@ III-C Depth/Mask Image Construction
$I_{M}(u,v)=\begin{cases}1&\text{if }I(u,v)\neq\rho_{\rm null},\\ 0&\text{otherwise}.\end{cases}$ (3)
@@
$\mathcal{D}_{M}=\{((u,v),I_{M}(u,v))\mid\$$u\in\{1,\ldots,W\},v\in\{1,\ldots,H\}\}.$ (4)
@@
In addition, we divide the depth image into small rectangular areas, or patches, inspired by recent works [yunpengps-nerv] to improve the decoding performance and the quality of the reconstructed depth image. Specifically, the RI is evenly segmented into patches $I_{D}^{\prime}(i)\in\mathbb{R}^{\frac{W}{N_{p}}\times\frac{H}{N_{p}}}$, where $N_{p}$ is a scaling factor and $i=1,\cdots,N_{p}^{2}$. Each patch is assigned a patch index, allowing us to specify a pixel in the patched RI as $I_{D}^{\prime}(i,i_{u},i_{v})$, where $i$ represents the patch index and $i_{u},i_{v}$ denotes the in-patch pixel coordinates whose origin is the top left pixel in the $i$-th patch. Similarly to the mask image, we also construct a dataset $\mathcal{D}_{D}$ for training depth INR $\Psi(\cdot;\bm{\omega})$ which consists of pairs of a patch index, in-patch coordinates, and the corresponding depth values, excluding unassigned pixels, as follows:
$\mathcal{D}_{D}=\{((i,i_{u},i_{v}),I_{D}(i,i_{u},i_{v}))\mid\$$i\in\{1,\ldots,N_{p}^{2}\},$
@@ III-D INR-based RI Encoder
In the encoding process, the mask INR $\Phi(\cdot;\psi)$ and the depth INR $\Psi(\cdot;\omega)$ are trained to obtain good parameters to express the coordinate-to-value relationships contained in the mask dataset $\mathcal{D}_{M}$ and $\mathcal{D}_{D}$. Figs. 2 (a) and (b) show the detailed architecture of the proposed depth INR and mask INR, respectively.
@@ III-D2 Depth INR
Similar to the mask INR, we also assume the existence of another function $\Psi_{D}$ regarding the depth image, which maps the pair of a patch index and an in-patch pixel coordinate to a depth value as
$\Psi_{D}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{1}.$ (9)
and the objective of the depth INR is to obtain parameters $\bm{\omega}$ for good approximation of $\Psi_{D}$ as $\Psi((i,i_{u},i_{v});\bm{\omega})\approx\Psi_{D}$. Figs. 2 (b) shows the detailed architecture of the proposed depth INR. The depth INR shares its structure with the mask INR, with the exception of the input layer, which accepts 3 values. In the training process, pairs of patch index $i$ and an in-patch coordinate $(i_{u},i_{v})$ are sequentially passed to the depth INR network from the depth dataset $\mathcal{D}_{D}$, and the corresponding depth values are regressed. We employ mean squared error (MSE) loss as a regression loss function for the depth INR as
$\mathcal{L}_{\mathrm{MSE}}(\bm{\omega})=\frac{1}{HW}\sum_{i}^{N_{p}^{2}}\sum_{i_{u}}^{W/N_{p}}\sum_{i_{v}}^{H/N_{p}}\|\Psi((i,i_{u},i_{v});\bm{\omega})-I_{D}(i,i_{u},i_{v})\|^{2}.$ (10)